# Neural Solver Selection for Combinatorial Optimization

### v1

@@ A.10 Complete results with all neural solvers
Table 7: Empirical results on synthetic TSP and TSPLIB datasets, reporting the mean (standard deviation) over five independent runs. All individual solvers are included for comparison, and OPT denotes the optimal performance for selection, which is computed by running all individual solvers in the zoo for each instance and selecting the best one. The best individual solver and its results are underlined, and the best optimality gaps, excluding OPT, are highlighted in boldface. The suffixes ’-N100’ and ’-N500’ indicate models trained on datasets with N=100 and N=500, respectively.
@@
Gap Time
BQ 3.00% 1.61s
ELG 3.70% 0.45s
LEHD 3.57% 0.86s
T2T-N100 2.48% 1.71s
T2T-N500 2.40% 1.98s
DIFUSCO-N100 2.84% 1.46s
DIFUSCO-N500 2.33% 1.45s
OPT 1.24% 8.93s
@@
Gap Time
BQ 3.04% 1.44s
ELG 3.05% 0.40s
LEHD 2.57% 0.88s
T2T-N100 2.09% 1.76s
T2T-N500 1.95% 1.74s
DISFUCO-N100 2.25% 1.44s
DISFUCO-N500 2.13% 1.44s
OPT 0.89% 9.14s
@@
Greedy 1.33% (0.06%) 1.28s (0.03s)
Top-$k$ ($k=2$) 1.07% (0.03%) 2.48s (0.02s)
Rejection ($20\%$) 1.26% (0.03%) 1.51s (0.04s)
Top-$p$ ($p=0.5$) 1.28% (0.04%) 1.46s (0.06s)
Table 8: Empirical results on synthetic CVRP and CVRPLIB Set-X datasets, reporting the mean (standard deviation) over five independent runs. All individual solvers are included for comparison.
@@
Gap Time
BQ 7.20% 1.59s
ELG 7.81% 0.82s
LEHD 7.37% 1.01s
MVMoE 13.56% 0.70s
Omni 6.82% 0.24s
OPT 4.64% 4.38s
@@
Gap Time
BQ 10.31% 2.60s
ELG 6.10% 1.31s
LEHD 13.70% 1.60s
MVMoE 15.54% 0.90s
Omni 6.21% 0.38s
OPT 5.10% 6.81s

### v2

@@ A.12 Complete results with all neural solvers
Table 10: Empirical results on synthetic TSP and TSPLIB datasets, reporting the mean (standard deviation) over five independent runs. All individual solvers are included for comparison. The suffixes ’-N100’ and ’-N500’ indicate models trained on datasets with N=100 and N=500, respectively. Obj denotes the average objective value on the dataset.
@@
      Obj       Gap       Time
      BQ       8.13       3.00%       1.61s
      ELG       8.18       3.70%       0.45s
      LEHD       8.17       3.57%       0.86s
      T2T-N100       8.08       2.48%       1.71s
      T2T-N500       8.08       2.40%       1.98s
      DIFUSCO-N100       8.11       2.84%       1.46s
      DIFUSCO-N500       8.07       2.33%       1.45s
      Oracle       7.99       1.24%       8.93s
@@
      Obj       Gap       Time
      BQ       8.29       3.04%       1.44s
      ELG       8.29       3.05%       0.40s
      LEHD       8.26       2.57%       0.88s
      T2T-N100       8.22       2.09%       1.76s
      T2T-N500       8.21       1.95%       1.74s
      DIFUSCO-N100       8.23       2.25%       1.44s
      DIFUSCO-N500       8.22       2.13%       1.44s
      Oracle       8.12       0.89%       9.14s
@@
      Greedy       8.16 (0.00)       1.33% (0.06%)       1.28s (0.03s)
      Top-$k$ ($k=2$)       8.14 (0.00)       1.07% (0.03%)       2.48s (0.02s)
      Rejection ($20\%$)       8.15 (0.00)       1.26% (0.03%)       1.51s (0.04s)
      Top-$p$ ($p=0.5$)       8.15 (0.00)       1.28% (0.04%)       1.46s (0.06s)
Table 11: Empirical results on synthetic CVRP and CVRPLIB Set-X datasets, reporting the mean (standard deviation) over five independent runs. All individual solvers are included for comparison. Obj denotes the average objective value on the dataset.
@@
       Obj        Gap        Time
       BQ        18.39        7.20%        1.59s
       ELG        18.49        7.81%        0.82s
       LEHD        18.42        7.37%        1.01s
       MVMoE        19.48        13.56%        0.70s
       Omni        18.32        6.82%        0.24s
       Oracle        17.95        4.64%        4.38s
@@
       Obj        Gap        Time
       BQ        71.21        10.31%        2.60s
       ELG        68.50        6.10%        1.31s
       LEHD        73.40        13.70%        1.60s
       MVMoE        74.59        15.54%        0.90s
       Omni        68.57        6.21%        0.38s
       Oracle        67.85        5.10%        6.81s